Complement Probability Rule

Rarely Tested

Complement Probability Rule

For any event A:

$$P(A^c)=1-P(A)$$

Usage

- Used when calculating the probability of the complement is easier.

Question 1

A box contains $$n$$ identical blue balls and $$3$$ identical red balls. Two balls are drawn at random without replacement. If the probability that both drawn balls are blue is exactly equal to the probability that they are of different colors, let this specific value of $$n$$ be $$N$$. If $$N$$ blue balls and $$3$$ red balls are now placed back into the box, and balls are drawn one by one with replacement, the probability that at most 3 draws are required to obtain a red ball is:

Go back to topics

Join CAT 2026 course by 5-Time CAT 100%iler

Start your IIM journey with the right preparation and crack CAT 2026.